Two concentic rings placed in a gravity free free region in yz plane of radius R carries a charge of +Q and second of Radius 4R and charge -8Q distributed uniformly over it. Find the minimum velocity with which a point charge of mass m and charge -q should be projected from a point at a distance 3R from the centre of the rings on its axis so that it will reach to the centre of the rings.

Grade:12

1 Answers

Sher Mohammad IIT Delhi
9 years ago
*force on charge q is = F1-F2, F1 is repulsive force and F2 is attractive force due to negative charge.
distance is s=3r
s=ut-1/2*a*t^2
s+.5at^2=ut
or 0=u-at
sher mohammad
faculty askiitians
b.tech, iit delhi